OpenStudy (anonymous):

if the 4th term of an arithmetic sequence is 31 and the 8th term of the sequence is 59, what is the first term of the sequence?

OpenStudy (anonymous):

4th term = a + 3d = 31 8th term = a + 7d = 59 after solving d= 7, a = 10 so, first term = 10, and sequence is : 10, 17, 24, 31,........